OpenStudy (anonymous):

Line PB has endpoints at -2 and 12. What is the coordinate of it midpoint??????

OpenStudy (anonymous):

10?

OpenStudy (anonymous):

5

OpenStudy (anonymous):

14

OpenStudy (gorv):

12-2=10/2=5 you got right

OpenStudy (anonymous):

oh okay